OpenStudy (anonymous):

someone help me solve these two questions 1) Prove that Lim x-->-1- 5/(x+1)^3=-infinity 2)prove using definition 8, that lim x->infinity x^3=infinity

OpenStudy (anonymous):

let L= \[\lim_{x \rightarrow -1-}\frac{ -5 }{ \left( x+1 \right)^3 }\] put x=-1-h,h>0, \[h \rightarrow0~as~x \rightarrow-1-\] \[L=\lim_{h \rightarrow 0}\frac{ 5 }{ \left( -1-h +1\right)^3 }=\lim_{h \rightarrow 0}\frac{ -5 }{ h^3 }\rightarrow -\infty \]

OpenStudy (anonymous):

Your definitions may vary, but here are the ones I'll be referring to: One-sided limits: \[\large\lim_{x\to c^-}f(x)=L\] if for any \(\epsilon>0\), there exists \(\delta>0\) such that \[\large-\delta<x-c<0~~\implies~~|f(x)-L|<\epsilon\] Infinite limits: \[\large\lim_{x\to c}f(x)=-\infty\] if for any \(N<0\) there exists \(\delta>0\) such that \[\large0<|x-c|<\delta~~\implies~~f(x)<N\] For your particular limit, \[\large\lim_{x\to-1^-}\frac{5}{(x+1)^3}=-\infty\] you have to show that for any \(M<0\) there is some \(\delta>0\) such that \[\large-\delta<|x+1|<0~~\implies~~\frac{5}{(x+1)^3}<M\]

OpenStudy (anonymous):

Oops, that last line should read \[\large -\delta<x+1<0~~\implies\cdots\]